Evaluate: |–11 – 4| – |1 – 5|: Evaluate: |–11 – 4| – |1 – 5|
Answer by mangopeeler07(445)

|–11 – 4| – |1 – 5|

-11-4=-15
absolute value:15

1-5=-4
absolute value:4

15-4=11

Answer: 11